# Poincare field theory for massless particles

### v2

@@ 12 Conclusion
Our basic assumption is that for arbitrary symmetry group all field equations follow from corresponding Casimir eigenvalue equations. In the present article we confirmed that statement for the case of massless Poincare group.
Our systematic approach has verified Weinberg observation [4] that Lorentz invariance is violated with terms in the form of local gauge transformations. The source of this effect is non commutativity of components of Pauli-Lubanski vector. We check this statements for tensor fields with arbitrary rank.
Let us stress that helicity is crucial for determining field equations for irreducible components of some field $\Psi^{A}$ and to find the number of theirs degrees of freedom. Note that we can describe the same theory using different fields. For example, as it shows Eq.(8) all representations with helicities $\lambda=\pm 1$ (vector field, symmetric and antisymmetric second rank tensor) have one more vector $\breve{p}^{a}_{\pm}$ than $\breve{p}^{a}_{\mp}$ and describe electromagnetic interaction. Also, all representations with helicities $\lambda=\pm 2$ (symmetric second rank tensor and fourth rank tensor with appropriate symmetries) have two more vector $\breve{p}^{a}_{\pm}$ than $\breve{p}^{a}_{\mp}$ and describe gravitation interaction. Consequently, contrary to the common belief that only highes helicity of certain tensor is relevant, we claim that all helicities are relevant. Different fields with the same helicity can describe the same theory, but sometimes it not easy to write field equation in terms of that fields.
Although the main contribution of the present article is proof that principle field equations (8) produce all known free field theory equations, this paper is also improvement over the set of articles with highest helicity [5, 6, 7, 8], well known in literature. For details and relations with these papers see Section $11$.
@@
It should be emphasize the important logical sequence in the present article Covariant definition of helicity $\to$ little group invariance $\to$ local gauge transformations $\to$ equations of motion $\to$ actions for massless fields. In other words, the necessity of existence of local gauge transformations for massless fields originates from requirement for Lorentz invariance.

### v3

@@ 13 Conclusion
Our fundamental assumption is that all field equations for an arbitrary symmetry group can be derived from the corresponding Casimir eigenvalue equations. In this work, we explicitly confirm this statement for the case of the massless Poincaré group.
Through our systematic approach, we have verified Weinberg’s observation [4] that Lorentz invariance is violated by terms taking the form of local gauge transformations. The principle field equations provide a simpler confirmation of Weinberg’s approach by working directly with fields rather than creation and annihilation operators. Indeed, these equations allow for a more straightforward derivation of the corresponding actions, which play a crucial role in quantization. The origin of Weinberg’s observation lies in the non-commutativity of the Pauli-Lubanski vector components. We have rigorously verified these results for tensor fields of arbitrary rank.
The introduction of the projectors is crucial because their diagonalization of the characteristic polynomial allows us to solve the corresponding determinants of arbitrary order and determine the spectrum of all helicities.
We defined the basis vectors such that the helicity $\lambda$ takes a simple form in terms of these vectors. Specifically, we have $\lambda=n_{+}-n_{-}$, where $n_{\pm}$ denotes the number of basis vectors $\breve{p}_{\pm}^{a}$. We emphasize that helicity plays a fundamental role in determining both the field equations for irreducible components of a given field $\Psi^{A}$ and the number of their physical degrees of freedom. Importantly, the same physical theory can be described using different field representations. For instance, as demonstrated in Eq.(8) all representations with helicities $\lambda=\pm 1$ (including vector fields, symmetric and antisymmetric second-rank tensors and arbitrary rank tensors ) exhibit one additional vector $\breve{p}^{a}_{\pm}$ compared to $\breve{p}^{a}_{\mp}$ and consistently describe electromagnetic interactions.
Similarly, representations with helicities $\lambda=\pm 2$ (such as symmetric second-rank tensors and fourth-rank tensors with appropriate symmetry properties) possess two additional vectors $\breve{p}^{a}_{\pm}$ over $\breve{p}^{a}_{\pm}$ and naturally describe gravitation interaction.
These results challenge the conventional view that only the highest helicity of a given tensor field is physically significant. Instead, we demonstrate that all helicity states are physically relevant. While different field representations with the same helicity can describe identical physics, formulating the field equations in terms of certain representations may prove non-trivial in practice.
While the primary contribution of this work is the demonstration that the principle field equations (8) generate all known free field theory equations, our results also represent a significant advancement beyond previous studies of highest helicity fields [5, 6, 7, 8]. In all existing literature, only free relativistic fields with the highest helicity, based to Fierz-Pauli constraints, have been studied. This represents a particular case within present approach, with totally symmetric tensor fields. Unlike previous studies limited to highest-helicity fields, this approach also covers lower-helicity not totally symmetric fields. These fields admit a clear physical interpretation, as detailed in Sections 7, 8, and 12.
@@
A crucial logical progression emerges in our framework: Covariant helicity definition $\to$ little group invariance $\to$ local gauge transformations $\to$ equations of motion $\to$ actions for massless fields.
This sequence reveals that the requirement of Lorentz invariance necessarily leads to the existence of local gauge transformations for massless fields.